Mike Dooley's regular hourly wage is $16 an hour. He receives overtime pay at the rate of time and a half. The FICA tax rate is 7.65%. Mike is paid weekly. For the first pay period in January, Mike worked 49 hours. All hours over 40 earn the overtime rate. Mike's federal income tax withholding is $150 and his state income tax withholding is $60. Mike has authorized that $25 be withheld from his check each pay period for savings bonds.

Compute Mike Dooley's net pay for the pay period. (Round answers to 2 decimal places, e.g. 815.25)
Answer #1

Gross pay = (40*16) + (9*16*1.5) = 856

Net pay = Gross pay - deductions

= 856 - 7.65% - 150 - 60 - 25

= 555.52

Comment if you face any issues